# An operator learning perspective on parameter-to-observable maps

### v1

@@ 4.1.2. Assumptions
We work under three primary assumptions that involve the data, the prior, and the truth. The first assumption concerns the covariance operators that characterize the end-to-end learning framework (EE).
@@
    (simultaneous diagonalization) The covariance operator $\Sigma=\operatorname{Cov}(\nu)$ of input training data distribution $\nu$ satisfies (11). The covariance operators $\Sigma^{\prime}\coloneqq\operatorname{Cov}(\nu^{\prime})$ of the input test data distribution $\nu^{\prime}$ and $\Lambda$ of the prior are diagonalized in the eigenbasis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ of $\Sigma$ and have the representations
    $\Sigma^{\prime}=\sum_{j=1}^{\infty}\sigma_{j}^{\prime}\varphi_{j}\otimes\varphi_{j}\quad\text{and}\quad\Lambda=\sum_{j=1}^{\infty}\lambda_{j}\varphi_{j}\otimes\varphi_{j}\,.$ (19)
@@
Although in traditional linear inverse problems the simultaneous diagonalizability (A1) of the normal operator and the prior covariance is considered a strong assumption [2, 62], in the linear functional regression setting here we interpret this condition more mildly. Indeed, $\Sigma$ and $\Lambda$ only commute in the infinite data limit. The actual normal operator corresponding to the inverse problem (17) is $S_{N}^{*}S_{N}/N=\widehat{\Sigma}$ which does not commute with $\Lambda$ in general for any finite $N$. Moreover, it is often the case—especially in operator learning—that the data generation procedure is controlled by the practitioner. In this case, it is simple to choose the data and prior covariance operators to have the same eigenfunctions. For example, Matérn-like Gaussian measures with different regularity exponents and lengthscales are a common choice for the data measure in operator learning and for the prior measure in Bayesian inverse problems; it is natural to assume their covariances share the same eigenbasis. However, the assumption that $\Sigma^{\prime}$ and $\Lambda$ have the same eigenbasis is stronger because the test distribution $\nu^{\prime}$ may be outside of the user’s control and differ substantially from the training distribution in some applications. Finally, the power law eigenvalue decay conditions (A2) and (A3) are standard in learning theory and help to facilitate explicit convergence rates.
We need a relatively strong assumption about the Karhunen–Loève (KL) expansion of the training data distribution $\nu$ to obtain high probability error bounds.
@@
The requirement of independence of the KL coefficients, while commonly found in the literature [4, 26], is undesirable. Indeed, it rules out the special setting of kernel ridge regression of nonlinear functions. The independence of the KL expansion coefficients of $\nu$ is only used in the proof of Lemma C.20. This lemma requires a strong notion of subgaussian (108); the assumed independence of the KL coefficients suffices to satisfy this condition. It is of interest to determine other sufficient conditions compatible with (108) that do not require independence.
Last, we require a regularity assumption on the true linear functional $f^{\dagger}$ in order to derive convergence rates.
@@ 4.2. Full-field learning
In this subsection, we provide the additional assumptions and framework required for the setting in which the true linear PtO map $f^{\dagger}$ is factorized as $f^{\dagger}=q^{\dagger}\circ L^{\dagger}$, where the QoI $q^{\dagger}$ is a linear functional on $H$ and $L^{\dagger}$ is a self-adjoint linear operator on $H$. We allow $q^{\dagger}$ or $L^{\dagger}$ to potentially be unbounded with respect to the topology of $H$. With the full-field learning data access model (FF), we fully observe noisy versions of the function-valued output of $L^{\dagger}$ at the training input functions. We adopt a Bayesian posterior mean estimator $\bar{L}^{(N)}$ for $L^{\dagger}$ based on these data. The final estimator of the true linear functional $f^{\dagger}$ is obtained by composing $q^{\dagger}$ with the learned operator $\bar{L}^{(N)}$. Subsection 4.2.1 contains the setup, while Subsection 4.2.2 contains the assumptions we invoke and gives examples of some QoIs that satisfy these assumptions.
@@ 4.2.1. Setup and estimator
$Y_{n}=L^{\dagger}u_{n}+\eta_{n}\,,\quad\text{where}\quad u_{n}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\nu\quad\text{and}\quad\eta_{n}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,\operatorname{Id}_{H})$ (25)
for $n\in[N]$. Equation (25) should be interpreted in a weak sense, i.e., as $H$-indexed stochastic processes, because $\eta_{n}\not\in H$ a.s. [16, Section 2.2.2., p. 11]. Without loss of generality, we assume a unit noise level because this does not affect the asymptotic results. To make the analysis tractable, we work in the setting that $L^{\dagger}$ is diagonalized in the eigenbasis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ of $\Sigma$ from (11). Thus, we write
@@
Our Bayesian approach follows [16] by modeling the eigenvalue sequence $l^{\dagger}$ with an independent Gaussian prior $l_{j}\sim\mathcal{N}(0,\mu_{j})$ on each eigenvalue. Write $\Upsilon=\{Y_{n}\}_{n=1}^{N}$ and $U=\{u_{n}\}_{n=1}^{N}$ and assume that $l=\{l_{j}\}_{n\in\mathbb{N}}$, $\Upsilon$, and $U$ are independent. Then [16, Fact 2.4, p. 12] furnishes the posterior distribution
@@
In (27), $\{\bar{l}_{j}^{(N)}\}_{j\in\mathbb{N}}$ are the posterior mean eigenvalues and $\{c_{j}^{(N)}\}_{j\in\mathbb{N}}$ the posterior variances. The plug-in estimator for $f^{\dagger}=q^{\dagger}\circ L^{\dagger}$ is then given by
@@
The precise formulas for the mean and variance in (27) are given in [16, Equation (2.4), p. 12]. As in Subsection 4.1, we are interested in the out-of-distribution test squared error of $q^{\dagger}\circ\bar{L}^{(N)}$ with respect to input test measure $\nu^{\prime}$.
@@ 4.2.2. Assumptions
We now collect the main assumptions for the full-field learning approach; these are primarily drawn from [16, Assumption 3.1, pp. 14–15].
@@
    The input distributions $\nu$ and $\nu^{\prime}$ and their covariance operators $\Sigma$ and $\Sigma^{\prime}$ satisfy Assumptions (A1) and (A2). [^6] Inspection of the proof of [16, Thm. 3.9, pp. 18–19] shows that the high probability upper bound (Equation 3.10) there remains valid if it is assumed that $\sigma^{\prime}_{j}$ is only bounded above asymptotically by $j^{-2\alpha^{\prime}}$ and not necessarily from below. Moreover, $\nu$ satisfies Assumption 4.3.
@@
    The eigenvalues $l^{\dagger}$ of the operator $L^{\dagger}$ in (26) satisfy $l^{\dagger}\in\mathcal{H}^{\beta}$ for some $\beta\in\mathbb{R}$.
@@
    The prior $l_{j}\sim\mathcal{N}(0,\mu_{j})$ has variances satisfying $\mu_{j}\asymp j^{-2\beta-1}$ as $j\to\infty$.
@@
    The QoI $q^{\dagger}$ satisfies $\lvert q^{\dagger}(\varphi_{j})\rvert^{2}\lesssim j^{-2r-1}$ as $j\to\infty$ for some $r\in\mathbb{R}$ such that $\min(\alpha,\alpha^{\prime}+r+1/2)+\beta>0$, where $\alpha$ and $\alpha^{\prime}$ are as in (A-I).
@@
    The PtO map $q^{\dagger}\circ L^{\dagger}$ is continuous: $\sum_{j=1}^{\infty}\lvert q^{\dagger}(\varphi_{j})\rvert^{2}\lvert l^{\dagger}_{j}\rvert^{2}<\infty$.
The conditions in Assumption 4.5 have similar interpretations to those in Subsection 4.1.2 and [16, Assumption 3.1, pp. 14–15]. For simplicity, in (A-III) we have already chosen the optimal prior smoothness exponent $\beta+1/2$. The condition $\alpha^{\prime}+r+\beta>0$ in (A-IV) ensures that the PtO map $q^{\dagger}\circ L^{\dagger}$ has finite $L^{2}_{\nu^{\prime}}(H;\mathbb{R})$ Bochner norm; thus, the test error is well-defined. The continuity of the PtO map enforced by (A-V) aligns with the (EE) setting. The power law decay of the coefficients of $q^{\dagger}$ in (A-IV) allows for a sharp convergence analysis. Several common, concrete linear QoIs satisfy this condition, as the next remark demonstrates.
@@
Several simple QoIs $q^{\dagger}$ satisfy the power law decay in (A-IV). Let $H=L^{2}((0,1);\mathbb{R})$, which has orthonormal basis $x\mapsto\varphi_{j}(x)=\sqrt{2}\sin(j\pi x)$ for each $j\in\mathbb{N}$. For convenience, denote $q^{\dagger}_{j}\coloneqq q^{\dagger}(\varphi_{j})$.
@@
    $\lvert q^{\dagger}_{j}\rvert^{2}=\bigg{\lvert}\frac{\sqrt{2}(1-\cos(j\pi))}{j\pi}\bigg{\rvert}^{2}=\frac{8\mathbbm{1}_{\{j\ \text{odd}\}}}{j^{2}\pi^{2}}\lesssim j^{-2}\,.$
@@
    $\lvert q^{\dagger}_{j}\rvert^{2}=2\lvert\sin(2\pi jx_{0})\rvert^{2}\lesssim 1$
@@
    $\lvert q^{\dagger}_{j}\rvert^{2}=8\pi^{2}j^{2}\lvert\cos(2\pi jx_{0})\rvert^{2}\lesssim j^{2}$
@@ 4.3. Main results
Building off of the setup from the previous two subsections, this subsection establishes convergence rates for end-to-end learning in Theorem 4.7 and full-field learning in Theorem 4.8. Both results are stated for the expectation of the out-of-distribution test error (18) conditioned on the input data $U$. Intuitively, this averages out the noise in the data. The two theorems are interpreted in Subsection 4.3.1. This discussion is followed by Subsection 4.3.2, which directly compares the end-to-end and full-field methods in Corollary 4.10.
@@
Let the input
training data distribution $\nu$, the test data distribution $\nu^{\prime}$, and the Gaussian prior $\mathcal{N}(0,\Lambda)$ satisfy Assumptions 4.2 and 4.3. Let the ground truth linear functional $f^{\dagger}\in\mathcal{H}^{s}$ satisfy Assumption 4.4 with $s>0$. Let $\alpha$ and $\alpha^{\prime}$ be as in (20) and $p=s+1/2$ be as in (21). Then there exists $c\in(0,1/4)$ and $N_{0}\geq 1$ such that for any $N\geq N_{0}$, the mean $\bar{f}^{(N)}$ of the Gaussian posterior distribution (15) arising from the $N$ pairs of observed training data $(U,Y)$ in (17) satisfies the error bound
@@
The second main theorem in this subsection describes the expected squared error in the QoI after learning an approximate forward map from full-field data.

### v2

@@ 4.1.2. Assumptions
In the end-to-end setting, we work under three primary assumptions that involve the data, the prior, and the truth. The first assumption concerns the covariance operators that characterize the end-to-end learning framework (EE).
@@
    (simultaneous diagonalization) The covariance operator $\Sigma=\operatorname{Cov}(\nu)$ of centered input training data distribution $\nu$ satisfies (11). The covariance operators $\Sigma^{\prime}\coloneqq\operatorname{Cov}(\nu^{\prime})$ of the centered input test data distribution $\nu^{\prime}$ and $\Lambda$ of the centered Gaussian prior are diagonalized in the $H$-orthonormal eigenbasis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ of $\Sigma$ and have the representations
    $\Sigma^{\prime}=\sum_{j=1}^{\infty}\sigma_{j}^{\prime}\varphi_{j}\otimes\varphi_{j}\quad\text{and}\quad\Lambda=\sum_{j=1}^{\infty}\lambda_{j}\varphi_{j}\otimes\varphi_{j}$ (19)
    for some sequences $\{\sigma_{j}^{\prime}\}_{j\in\mathbb{N}}\subset\mathbb{R}_{\geq 0}$ and $\{\lambda_{j}\}_{j\in\mathbb{N}}\subset\mathbb{R}_{\geq 0}$.
@@
The power law eigenvalue decay conditions (A2) and (A3) are standard in learning theory and help to facilitate explicit algebraic convergence rates. However, other eigenvalue assumptions such as exponential decay [4] or convex eigenvalues [13] are also possible and interesting to study. The exponent lower bounds $\alpha>1/2$ and $p>1/2$ ensure that $\Sigma$ and $\Lambda$ are trace-class on $H$, and thus the posterior (16) is a proper Gaussian measure on $H$.
The simultaneous diagonalization condition (A1) is the strongest assumption in the paper and is the main limitation of our theory in Section 4. It implies that the prior covariance operator $\Lambda$, which we choose a priori, is perfectly aligned with the eigenspaces of the training and test data covariance operators $\Sigma$ and $\Sigma^{\prime}$, which are unknown in general. Nonetheless, in certain settings, these eigenspaces may be known from problem-specific knowledge such as periodicity, translation invariance, or rotation invariance. Also, if the data generation procedure or experimental design is controlled by the practitioner, then the data and prior covariance operators can freely be chosen to share the same eigenfunctions, e.g., Matérn-like covariances with different lengthscales and regularity exponents.
In our analysis and in related work [17, 34], the simultaneous diagonalization implied by Assumption (A1) enables the derivation of tight error bounds that remain valid for misspecified Gaussian priors. The resulting convergence rates clearly isolate the influence that problem regularity, data quality, and prior information have on sample complexity. In contrast, papers in statistical inverse problems that do not require the assumption of simultaneous diagonalization usually obtain less precise and less interpretable results compared to ours [3, 25, 32, 73, 85, 89]. Moreover, many of these works are limited to the well-specified setting in which the implied smoothness of the prior perfectly matches the unknown regularity of the ground truth. Our analysis overcomes this limitation at the expense of simultaneous diagonalization. In further contrast to traditional linear inverse problems, in our finite data linear functional regression setting the actual normal operator corresponding to the inverse problem (17) is the empirical covariance $S_{N}^{*}S_{N}^{\phantom{*}}/N=\widehat{\Sigma}$ (14), which does not commute with $\Lambda$ in general for any finite $N$. Our theory accommodates this challenge and only assumes that $\Lambda$ commutes with $\Sigma$ (11), which is the infinite data limit of $\widehat{\Sigma}$. However, the assumption that test data covariance $\Sigma^{\prime}$ and prior covariance $\Lambda$ have the same eigenbasis is less realistic because the test distribution $\nu^{\prime}$ may be outside of the user’s control and differ substantially from the training distribution $\nu$ (11) in some applications. Nevertheless, the present work is only one of a handful [17, 31, 61] to provide out-of-distribution test error bounds at all. The commuting assumption is satisfied for two popular choices of test data covariance, $\Sigma^{\prime}=\Sigma$ (if $\Sigma$ and $\Lambda$ commute) and $\Sigma^{\prime}=\operatorname{Id}_{H}$, which reflect the prediction error and the strong estimation error, respectively. It is clear that weakening or removing Assumption (A1) while maintaining sharp, interpretable, and widely applicable theoretical insights for PtO map learning is an important direction for future research.
Assumption 4.2 provides fine-grained control of the second moments of the data and prior distributions. Next, we impose a slightly strong assumption about the tails of the Karhunen–Loève (KL) expansion of the training data distribution $\nu$ (11) in order to obtain high probability error bounds.
@@
Equation (23) in Assumption 4.3 implies that the training data distribution $\nu$ (11) is subgaussian in a relatively strong sense. This enables the use of exponential concentration inequalities in the proofs of forthcoming results. We further insist that the KL expansion coefficients $\{z_{j}\}_{j\in\mathbb{N}}$ (22) are an independent family in order to align with a similar assumption made for full-field learning in the next subsection. However, the following remark explains how this undesirable independence condition can be eliminated for end-to-end learning at the expense of worse tail bounds.
Remark 4.4 (independence of the KL coefficients).
The requirement of independence of the KL expansion coefficients $\{z_{j}\}_{j\in\mathbb{N}}$ (22), while commonly found in the literature [5, 30], is undesirable because it excludes many interesting statistical models. This condition is only used in Lemma C.20, which shows that a particular event has high probability. This lemma requires a strong notion of subgaussianity (113); the assumed independence of the KL coefficients suffices to satisfy this condition. However, we show in Lemma C.21 that the requirement of independence can be replaced by the condition that the $\{z_{j}\}_{j\in\mathbb{N}}$ are pairwise uncorrelated. At the cost of a smaller probability for the event of interest, this improvement extends the applicability of our end-to-end learning theory. For example, using the kernel trick, the main theoretical results of this section imply convergence rates for Gaussian process regression of scalar-valued nonlinear functions as a special case.
Last, we require a Sobolev-like regularity assumption on the true linear functional $f^{\dagger}$ in order to derive explicit convergence rates.
@@ 4.2. Full-field learning
In this subsection, we provide the additional assumptions and framework required for the setting in which the true linear PtO map $f^{\dagger}$ is factorized as $f^{\dagger}=q^{\dagger}\circ L^{\dagger}$, where the QoI $q^{\dagger}$ is a linear functional on $H$ and $L^{\dagger}$ is a self-adjoint linear operator mapping $H$ to $H$. We allow $q^{\dagger}$ or $L^{\dagger}$ to potentially be unbounded with respect to the topology of $H$. With the full-field learning data access model (FF), we fully observe noisy versions of the function-valued output of $L^{\dagger}$ at the training input functions drawn from $\nu$ (11). We adopt a Bayesian posterior mean estimator $\bar{L}^{(N)}$ for $L^{\dagger}$ based on these data. The final estimator of the true linear functional $f^{\dagger}$ is obtained by composing $q^{\dagger}$ with the learned operator $\bar{L}^{(N)}$. Subsection 4.2.1 contains the setup, while Subsection 4.2.2 contains the assumptions we make and gives examples of some QoIs that satisfy these assumptions.
@@ 4.2.1. Setup and estimator
$Y_{n}=L^{\dagger}u_{n}+\eta_{n}\,,\quad\text{where}\quad u_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\nu\quad\text{and}\quad\eta_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,\operatorname{Id}_{H})$ (25)
for $n\in[N]$. Equation (25) should be interpreted in a weak sense, i.e., as $H$-indexed stochastic processes, because $\eta_{n}\not\in H$ almost surely [17, Section 2.2.2., p. 11]. Without loss of generality, we assume a unit noise level because this does not affect the asymptotic results. To make the analysis tractable, we work in the setting that $L^{\dagger}$ is diagonalized in the eigenbasis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ of $\Sigma$ from (11). Thus, we write
@@
Our Bayesian approach follows [17] by modeling the eigenvalue sequence $l^{\dagger}$ with an independent scalar Gaussian prior $l_{j}\sim\mathcal{N}(0,\mu_{j})$ on each eigenvalue. Write $\Upsilon=\{Y_{n}\}_{n=1}^{N}$ and $U=\{u_{n}\}_{n=1}^{N}$ and assume that $l=\{l_{j}\}_{n\in\mathbb{N}}$, $\Upsilon$, and $U$ are independent. Then [17, Fact 2.4, p. 12] furnishes the posterior distribution
@@
In (27), $\{\bar{l}_{j}^{(N)}\}_{j\in\mathbb{N}}$ are the posterior mean eigenvalues and $\{c_{j}^{(N)}\}_{j\in\mathbb{N}}$ are the posterior variances. The plug-in estimator for $f^{\dagger}=q^{\dagger}\circ L^{\dagger}$ is then given by
@@
The precise formulas for the mean and variance in (27) are given in [17, Equation (2.4), p. 12]. As in Subsection 4.1.1, we are interested in the out-of-distribution test squared error of $q^{\dagger}\circ\bar{L}^{(N)}$ with respect to a centered input test measure $\nu^{\prime}$.
@@ 4.2.2. Assumptions
We now collect the main assumptions for the full-field learning approach; these are primarily drawn from [17, Assumption 3.1, pp. 14–15].
@@
    (simultaneous diagonalization, data decay, and tail condition) The input distributions $\nu$ and $\nu^{\prime}$ and their covariance operators $\Sigma$ and $\Sigma^{\prime}$ satisfy Assumptions (A1) and (A2). [^6] Inspection of the proof of [17, Theorem 3.9, pp. 18–19] shows that the high probability upper bound (Equation 3.10) there remains valid if it is assumed that $\sigma^{\prime}_{j}$ is only bounded above asymptotically by $j^{-2\alpha^{\prime}}$ and not necessarily from below. Moreover, $\nu$ satisfies Assumption 4.3.
@@
    (forward operator regularity) The eigenvalues $l^{\dagger}$ of operator $L^{\dagger}$ in (26) satisfy $l^{\dagger}\in\mathcal{H}^{\beta}$ for some $\beta\in\mathbb{R}$ [^7] Recall that the Sobolev-like sequence Hilbert spaces $\mathcal{H}^{s}$ for $s\in\mathbb{R}$ are defined in Subsection 1.3.
@@
    (prior regularity) The prior $l_{j}\sim\mathcal{N}(0,\mu_{j})$ on eigenvalues has variances satisfying $\mu_{j}\asymp j^{-2\beta-1}$ as $j\to\infty$.
@@
    (QoI regularity) The QoI $q^{\dagger}$ satisfies $\lvert q^{\dagger}(\varphi_{j})\rvert^{2}\lesssim j^{-2r-1}$ as $j\to\infty$ for some $r\in\mathbb{R}$ such that $\min(\alpha,\alpha^{\prime}+r+1/2)+\beta>0$, with $\alpha$ and $\alpha^{\prime}$ as in (A-I).
@@
    (PtO map continuity) The PtO map $q^{\dagger}\circ L^{\dagger}$ is continuous, i.e.,
    $\sum_{j=1}^{\infty}\lvert q^{\dagger}(\varphi_{j})\rvert^{2}\lvert l^{\dagger}_{j}\rvert^{2}<\infty\,.$ (29)
The conditions in Assumption 4.6 have similar interpretations to those in Subsection 4.1.2 and [17, Assumption 3.1, pp. 14–15]. For simplicity, in (A-III) we have already chosen the optimal prior smoothness exponent $\beta+1/2$. The condition $\alpha^{\prime}+r+1/2+\beta>0$ in (A-IV) ensures that the PtO map $q^{\dagger}\circ L^{\dagger}$ has finite $L^{2}_{\nu^{\prime}}(H;\mathbb{R})$ Bochner norm; thus, the test error is well-defined. The continuity of the PtO map enforced by (A-V) aligns with the (EE) setting. The power law decay of the coefficients of $q^{\dagger}$ in (A-IV) allows for a sharp convergence analysis. Several common, concrete linear QoIs satisfy this condition, as the next remark demonstrates.
@@
Several simple QoIs $q^{\dagger}$ satisfy the power law decay in (A-IV). Let $H=L^{2}((0,1);\mathbb{R})$. An orthonormal basis for $H$ is $x\mapsto\varphi_{j}(x)=\sqrt{2}\sin(j\pi x)$ for each $j\in\mathbb{N}$. For convenience, denote $q^{\dagger}_{j}\coloneqq q^{\dagger}(\varphi_{j})$.
@@
    $\lvert q^{\dagger}_{j}\rvert^{2}=\bigg{\lvert}\frac{\sqrt{2}(1-\cos(j\pi))}{j\pi}\bigg{\rvert}^{2}=\frac{8\mathbbm{1}_{\{j\ \text{odd}\}}}{j^{2}\pi^{2}}\lesssim j^{-2}\,.$ (30)
@@
    $\lvert q^{\dagger}_{j}\rvert^{2}=2\lvert\sin(j\pi x_{0})\rvert^{2}\lesssim 1$ (31)
@@
    $\lvert q^{\dagger}_{j}\rvert^{2}=2\pi^{2}j^{2}\lvert\cos(j\pi x_{0})\rvert^{2}\lesssim j^{2}$ (32)
@@ 4.3. Main results
Building upon the setup from the previous two subsections, this subsection establishes convergence rates for end-to-end learning in Theorem 4.8 and full-field learning in Theorem 4.9. Both results are stated for the expectation of the out-of-distribution test error conditioned on the input data $U$. Intuitively, this averages out the noise in the data; full expectation bounds are also possible to obtain, as in [62, Corollary D.2, p. 226]. The two theorems are interpreted in Subsection 4.3.1. This discussion is followed by Subsection 4.3.2, which directly compares the end-to-end and full-field methods in Corollary 4.11.
@@
Let the input training data distribution $\nu$, the test data distribution $\nu^{\prime}$, and the Gaussian prior $\mathcal{N}(0,\Lambda)$ satisfy Assumptions 4.2 and 4.3. Let the ground truth linear functional $f^{\dagger}\in\mathcal{H}^{s}$ satisfy Assumption 4.5 with $s>0$. Let $\alpha$ and $\alpha^{\prime}$ be as in (20) and $p=s+1/2$ be as in (21). Then there exists $c\in(0,1/4)$ and $N_{0}\geq 1$ such that for any $N\geq N_{0}$, the mean $\bar{f}^{(N)}$ of the Gaussian posterior distribution (15) arising from the $N$ pairs of observed training data $(U,Y)$ in (17) satisfies the error bound
@@
The second main theorem in this subsection describes the expected squared test error in the QoI after learning an approximate forward map from full-field data.